OpenStudy (anonymous):

(Definite Intergral) from 0 to 1 (4)/1+x^2 dx by using trig functions

OpenStudy (anonymous):

\[\int_0^1\frac{4}{1+x^2}dx\]

OpenStudy (anonymous):

that is it

OpenStudy (anonymous):

boom shaka laka

OpenStudy (anonymous):

\[4\int_0^1 \frac{1}{1+x^2}dx \implies 4\left[\tan^{-1}x\right]_0^1 \implies 4\left(\frac{\pi}{4}- 0\right)\]